OpenStudy (anonymous):

Find the simplified form of the expression. Give your answer in scientific notation.

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

(3*10 -6 is above the 10)(6*10 -8 is above the 10)

OpenStudy (zehanz):

\(\dfrac{3 \cdot 10^{-6}}{6 \cdot 10^{-8}}\)

OpenStudy (zehanz):

You can split it up into \(\dfrac{3}{6} \cdot \dfrac{10^{-6}}{10^{-8}}\)

OpenStudy (zehanz):

You know what the outcome of 3/6 is. The other part: you have to divide two powers of 10. You can apply the rule: \(\dfrac{a^b}{a^c}=a^{b-c}\) So this would give you: \(0.5 \cdot 10^{-6-(-8)}\). Try to simplify.

OpenStudy (zehanz):

If you found that is is \(0.5 \cdot 10^{2}\), you're almost done! Remember, in scientific notation: \(a \cdot 10^{b}\), \(1 \le a<10\).